A brownie pan has dimensions
8" by 4" by 2 3/5".
Compute the volume of the pan.
Express your answer in decimal form.
If the batter is 1.75" deep, what percent of the pan is filled with brownie batter?
7th grade math

Answer by checkley75(3666)

8*4*2.6=83.2 CUBIC INCHES FOR THE VOLUME OF THIS PAN.
8*4*1.75=56 CUBIC INCHES FOR THE BATTER.
THUS 56/83.2=.673 OR 67.3% OF THE PAN IS FILLED WITH BATTER.